OpenStudy (anonymous):

h = 2 ft, w = 5 ft, l = 7 ft 28 ft3 70 ft3 90 ft3 118 ft3

OpenStudy (anonymous):

Is it a rectangle? Are you finding the area?

OpenStudy (anonymous):

Is there a picture?

OpenStudy (anonymous):

What is the volume of a rectangular prism with the following dimensions? Sorry forgot that

OpenStudy (anonymous):

Okay, here's the formula: \(A=2(wl+hl+hw)\) Plug in what we know: \(A=2((5)(7)+(2)(7)+(2)(5))\) Okay, can you multiply 5 * 7?

OpenStudy (anonymous):

35

OpenStudy (anonymous):

Yes, so we have: \(A=2(35+(2)(7)+(2)(5))\) Okay, can you multiply 2 * 7?

OpenStudy (anonymous):

14

OpenStudy (anonymous):

Yes, so we have: \(A=2(35+14+(2)(5))\) Okay, can you multiply 2 * 5?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

Yes, so we have: \(A=2(35+14+10)\) Okay, can you add 35 + 14 + 10?

OpenStudy (anonymous):

59

OpenStudy (anonymous):

Yes, so we have: \(A = 2(59)\) Can you multiply 2 * 59?

OpenStudy (anonymous):

118

OpenStudy (anonymous):

Yep, that's our answer!